OpenStudy (stuck-help):

what is the simplified for of the expression in the comments. answer in scientific notation need to show work

OpenStudy (stuck-help):

\[\left( 8\times10^{7} \right)\left( 7\times10^{4} \right)\]

OpenStudy (anonymous):

\[8\times 7\times 10^7\times 10^4\] is a start then add up the exponents, and also multiply \(8\times 7\) there is one more step after that

OpenStudy (stuck-help):

\[56\times10^{11}\]

OpenStudy (anonymous):

yeah good but still one more step

OpenStudy (stuck-help):

divid by 10 ?

OpenStudy (anonymous):

it is not in scientific notation, because \(56>10\) so turn it in to \(5.6\) then kick the exponent up by 1

OpenStudy (anonymous):

you should get as a final answer \[5.6\times 10^{12}\]